OpenStudy (anonymous):

log 10 x - log 10 (x-4) =1 ?

OpenStudy (anonymous):

\[\log_{10}x - \log_{10}x-4 =1\] \[\log_{10}\frac{x}{x-4} = \log_{10}10\] \[ \frac{x}{x-4} = 10\] \[x =10x - 40\] \[9x =40\] \[x =\frac{40}{9}\]